OpenStudy (anonymous):

how do I find my distance using the distance formula with points (-1,1) and (2,-3) also would the answer be 5 or the square root of 5?

OpenStudy (debbieg):

\(\Large d=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^2}\) Just plug in the x's and y's of your two points: (-1,1) and (2,-3) \(\Large d=\sqrt{(-1-2)^{2}+(1-(-3))^2}\) and simplify! What do you get for the answer?

OpenStudy (anonymous):

is it 5 or 5 squared?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

what did you get? the formula above is clear, isn't it?

OpenStudy (anonymous):

yes I got the answer 5 but is it s?

OpenStudy (anonymous):

5 squared?

OpenStudy (anonymous):

the distance is 5 units

OpenStudy (anonymous):

no, you took the sqrt root of 5 squared to get 5. that is the distance

OpenStudy (anonymous):

ohhh ok thank you

OpenStudy (anonymous):

you're welcome

OpenStudy (debbieg):

Do you understand why? How to get that from the formula? \(\Large d=\sqrt{(-1-2)^{2}+(1-(-3))^2}=\sqrt{(-3)^{2}+(4)^2}\\\Large =\sqrt{9+16}=?\)

OpenStudy (anonymous):

yes ty